Circular No. 8867

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 R1 (LARSON) S. Larson, Lunar and Planetary Laboratory, reports his discovery of

a comet on images taken in the course of the Mt. Lemmon Survey (discovery observation tabulated below), noting the object to have a 40′′ coma and a very diffuse tail in p.a. ∼ 230^{o}. Four co-added 60-s follow-up images by R. A. Kowalski with the Mt. Lemmon 1.5-m reflector on Sept. 8.4 UT in good seeing show a condensed comet 20′′ in diameter, slightly elongated toward p.a. ∼ 230^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, numerous other astrometric observers have also commented on the object’s cometary nature. G. Hug (Eskridge, KS, U.S.A., 0.7-m reflector) writes that his CCD frames taken on Sept. 8.3 show the object to be diffuse with a faint tail in p.a. ∼ 250^{o}. L. Buzzi and F. Luppi (Varese, Italy, 0.60-m f/4.64 reflector) report that their images from Sept. 8.89–9.07 show a strong central condensation surrounded by a coma at least 15′′ wide, with an extension towards the west. Fifteen stacked 60-s unfiltered exposures obtained during Sept. 9.03–9.08 by G. Sostero, E. Guido, and V. Gonano (Remanzacco, Italy, 0.45-m f/4.4 reflector) reveal the presence of a coma almost 12′′ in diameter. C. Jacques and E. Pimentel (Belo Horizonte, Minas Gerais, Brazil, 0.30-m f/3 Schmidt-Cassegrain reflector) find a coma diameter of 20′′ on their CCD frames from Sept. 9.1. E. Reina L. (Masquefa, Spain, 0.25-m f/3.3 Schmidt-Cassegrain reflector) finds the comet to be diffuse with a 23′′ coma on images taken on Sept. 9.07–9.15. J. Young (Table Mountain 0.61-m f/16 Cassegrain reflector) reports that his images from Sept. 9.24–9.27 show a 8′′ round coma with little central condensation and only a hint of some extension 12′′ long at p.a. 260^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Sept.4.37126  0^{h}10^{m}49.^{s}76  +1^{o}49′44′′.7  18.2

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2007-R36.

T = 2009 Feb. 22.484 TT  ω = 249.^{o}888
e = 0.40046  Ω = 181.461 }2000.0
q = 3.77499 AU  i = 9.725
a = 6.29652 AU n^{o} = 0.062381 P = 15.8 years
2007 September 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green